OpenStudy (anonymous):

A city's current population is 500,000 people. It is growing at a rate of 2.5% per year. The equation p = 500,000(1.025)^x models the city's population growth where x is the number of years from the current year. In approximately how many years will the population be 640,000?

OpenStudy (anonymous):

if p=640,000 you have: 640,000 = 500,000(1.025)^x solving for x: 640,000 / 500,000 = (1.025)^x using log in both sides of equation: log(640/500) = log(1.025)^x by applying properties you get: log(640/500) = x log(1.025) then: x = log(640/500)/log(1.025)

OpenStudy (anonymous):

sooo is the answer 15

OpenStudy (anonymous):

10